\begin{thebibliography}{10}

\bibitem{dorigo1999ant}
M~Dorigo and G~DiCaro.
\newblock The ant colony meta-heuristic in “new ideas in optimization”,
  1999.

\bibitem{Dorigo1999}
Marco Dorigo, Gianni~Di Caro, and Luca~M Gambardella.
\newblock Ant algorithms for discrete optimization.
\newblock {\em Artificial life}, 5(2):137--172, 1999.

\bibitem{dorigo1992optimization}
Marco Dorigo.
\newblock Optimization, learning and natural algorithms.
\newblock {\em PhD Thesis, Politecnico di Milano}, 1992.

\bibitem{dorigo1991ants}
Marco Dorigo, Vittorio Maniezzo, and Alberto Colorni.
\newblock The ant system: An autocatalytic optimizing process.
\newblock 1991.

\bibitem{dorigo1996ant69}
Marco Dorigo, Vittorio Maniezzo, and Alberto Colorni.
\newblock Ant system: optimization by a colony of cooperating agents.
\newblock {\em IEEE Transactions on Systems, Man, and Cybernetics, Part B
  (Cybernetics)}, 26(1):29--41, 1996.

\bibitem{applegate2006traveling6}
Robert~E Applegate, David L~Bixby, Vasek Chvatal, and William~J Cook.
\newblock {\em The traveling salesman problem: a computational study}.
\newblock Princeton university press, 2006.

\bibitem{reinelt1994traveling155}
Gerhard Reinelt.
\newblock {\em The traveling salesman: computational solutions for TSP
  applications}.
\newblock Springer-Verlag, 1994.

\bibitem{blum2008ant27}
Christian Blum, Mateu~Y{\'a}bar Vall{\`e}s, and Maria~J Blesa.
\newblock An ant colony optimization algorithm for dna sequencing by
  hybridization.
\newblock {\em Computers \& Operations Research}, 35(11):3620--3635, 2008.

\bibitem{blum2005beam20}
Christian Blum.
\newblock Beam-aco—hybridizing ant colony optimization with beam search: An
  application to open shop scheduling.
\newblock {\em Computers \& Operations Research}, 32(6):1565--1591, 2005.

\bibitem{korb2007ant107}
Oliver Korb, Thomas St{\"u}tzle, and Thomas~E Exner.
\newblock An ant colony optimization approach to flexible protein--ligand
  docking.
\newblock {\em Swarm Intelligence}, 1(2):115--134, 2007.

\bibitem{blum2008beam21}
Christian Blum.
\newblock Beam-aco for simple assembly line balancing.
\newblock {\em INFORMS Journal on Computing}, 20(4):618--627, 2008.

\bibitem{gambardella2000ant84}
Luca~Maria Gambardella and Marco Dorigo.
\newblock An ant colony system hybridized with a new local search for the
  sequential ordering problem.
\newblock {\em INFORMS Journal on Computing}, 12(3):237--255, 2000.

\bibitem{di1998antnet52}
Gianni Di~Caro and Marco Dorigo.
\newblock Antnet: Distributed stigmergetic control for communications networks.
\newblock {\em Journal of Artificial Intelligence Research}, 9:317--365, 1998.

\bibitem{shmygelska2005ant160}
Alena Shmygelska and Holger~H Hoos.
\newblock An ant colony optimisation algorithm for the 2d and 3d hydrophobic
  polar protein folding problem.
\newblock {\em BMC bioinformatics}, 6(1):30, 2005.

\bibitem{doerner2009special57}
Karl~F Doerner, Daniel Merkle, and Thomas St{\"u}tzle.
\newblock Special issue on ant colony optimization.
\newblock {\em Swarm Intelligence}, 3(1):1--2, 2009.

\bibitem{deneubourg1990self48}
J-L Deneubourg, Serge Aron, Simon Goss, and Jacques~M Pasteels.
\newblock The self-organizing exploratory pattern of the argentine ant.
\newblock {\em Journal of insect behavior}, 3(2):159--168, 1990.

\bibitem{beckers1993modulation10}
Ralph Beckers, Jean-Louis Deneubourg, and Simon Goss.
\newblock Modulation of trail laying in the antlasius niger (hymenoptera:
  Formicidae) and its role in the collective selection of a food source.
\newblock {\em Journal of Insect Behavior}, 6(6):751--759, 1993.

\bibitem{gambardella1995ant82}
Luca~M Gambardella and Marco Dorigo.
\newblock Ant-q: A reinforcement learning approach to the traveling salesman
  problem.
\newblock In {\em Machine Learning Proceedings 1995}, pages 252--260. Elsevier,
  1995.

\bibitem{dorrigo1997ant64}
M~Dorrigo and Luca~Maria Gambardella.
\newblock Ant colonies for the traveling salesman problem.
\newblock {\em in BioSystems}, 1997.

\bibitem{stutzle1996improving174}
Thomas St{\"u}tzle and Holger~H Hoos.
\newblock Improving the ant system: A detailed report on the max--min ant
  system.
\newblock {\em FG Intellektik, FB Informatik, TU Darmstadt, Germany, Tech. Rep.
  AIDA--96--12}, 1996.

\bibitem{stutzle1997max175}
Thomas Stutzle and Holger Hoos.
\newblock Max-min ant system and local search for the traveling salesman
  problem.
\newblock In {\em Evolutionary Computation, 1997., IEEE International
  Conference on}, pages 309--314. IEEE, 1997.

\bibitem{bullnheimer1997new31}
Bernd Bullnheimer, Richard~F Hartl, and Christine Strauss.
\newblock A new rank based version of the ant system. a computational study.
\newblock 1997.

\bibitem{maniezzo1999exact124}
Vittorio Maniezzo.
\newblock Exact and approximate nondeterministic tree-search procedures for the
  quadratic assignment problem.
\newblock {\em INFORMS journal on computing}, 11(4):358--369, 1999.

\bibitem{cordon2002analysis38}
Oscar Cord{\'o}n~Garc{\'\i}a, I{\~n}aki Fern{\'a}ndez~de Viana, and Francisco
  Herrera~Triguero.
\newblock Analysis of the best-worst ant system and its variants on the tsp.
\newblock {\em Mathware \& soft computing. 2002 Vol. 9 N{\'u}m. 2 [-3]}, 2002.

\bibitem{cordon2000new39}
Oscar Cordon, Inaki~Fern{\'a}ndez de~Viana, Francisco Herrera, and Llanos
  Moreno.
\newblock A new aco model integrating evolutionary computation concepts: The
  best-worst ant system.
\newblock 2000.

\bibitem{guntsch2002population92}
Michael Guntsch and Martin Middendorf.
\newblock A population based approach for aco.
\newblock In {\em Workshops on Applications of Evolutionary Computation}, pages
  72--81. Springer, 2002.

\bibitem{blum2004theoretical19}
Christian Blum.
\newblock {\em Theoretical and practical aspects of ant colony optimization},
  volume 282.
\newblock IOS press, 2004.

\bibitem{blum2001hcHCACO}
Christian Blum, Andrea Roli, and Marco Dorigo.
\newblock Hc--aco: The hyper-cube framework for ant colony optimization.
\newblock In {\em Proceedings of MIC}, volume~2, pages 399--403, 2001.

\bibitem{mavrovouniotis2017ant133}
Michalis Mavrovouniotis, Felipe~M M{\"u}ller, and Shengxiang Yang.
\newblock Ant colony optimization with local search for dynamic traveling
  salesman problems.
\newblock {\em IEEE transactions on cybernetics}, 47(7):1743--1756, 2017.

\bibitem{mavrovouniotis2015ant131}
Michalis Mavrovouniotis and Shengxiang Yang.
\newblock Ant algorithms with immigrants schemes for the dynamic vehicle
  routing problem.
\newblock {\em Information Sciences}, 294:456--477, 2015.

\bibitem{mavrovouniotis2017survey132}
Michalis Mavrovouniotis, Changhe Li, and Shengxiang Yang.
\newblock A survey of swarm intelligence for dynamic optimization: Algorithms
  and applications.
\newblock {\em Swarm and Evolutionary Computation}, 33:1--17, 2017.

\bibitem{van2013ant85}
Barry van Veen, Michael Emmerich, Zhiwei Yang, Thomas B{\"a}ck, and Joost Kok.
\newblock Ant colony algorithms for the dynamic vehicle routing problem with
  time windows.
\newblock In {\em International Work-Conference on the Interplay Between
  Natural and Artificial Computation}, pages 1--10. Springer, 2013.

\bibitem{lopez2012experimental119}
Manuel L{\'o}pez-Ib{\'a}{\~n}ez and Thomas St{\"u}tzle.
\newblock An experimental analysis of design choices of multi-objective ant
  colony optimization algorithms.
\newblock {\em Swarm Intelligence}, 6(3):207--232, 2012.

\bibitem{lopez2012automatic118}
Manuel Lopez-Ibanez and Thomas Stutzle.
\newblock The automatic design of multiobjective ant colony optimization
  algorithms.
\newblock {\em IEEE Transactions on Evolutionary Computation}, 16(6):861--875,
  2012.

\bibitem{falcon2017new77}
Jes{\'u}s~Guillermo Falc{\'o}n-Cardona and Carlos A~Coello Coello.
\newblock A new indicator-based many-objective ant colony optimizer for
  continuous search spaces.
\newblock {\em Swarm Intelligence}, 11(1):71--100, 2017.

\bibitem{liao2011incremental113}
Tianjun Liao, Marco~A Montes~de Oca, Dogan Aydin, Thomas St{\"u}tzle, and Marco
  Dorigo.
\newblock An incremental ant colony algorithm with local search for continuous
  optimization.
\newblock In {\em Proceedings of the 13th annual conference on Genetic and
  evolutionary computation}, pages 125--132. ACM, 2011.

\bibitem{kumar2015enhancing109}
Udit Kumar, Sumit Soman, et~al.
\newblock Enhancing iaco r local search by mtsls1-bfgs for continuous global
  optimization.
\newblock In {\em Proceedings of the 2015 Annual Conference on Genetic and
  Evolutionary Computation}, pages 33--40. ACM, 2015.

\bibitem{guo2012antNew}
Ping Guo and Lin Zhu.
\newblock Ant colony optimization for continuous domains.
\newblock In {\em Natural computation (ICNC), 2012 eighth international
  conference on}, pages 758--762. IEEE, 2012.

\bibitem{liao2014unified114}
Tianjun Liao, Thomas St{\"u}tzle, Marco A~Montes de~Oca, and Marco Dorigo.
\newblock A unified ant colony optimization algorithm for continuous
  optimization.
\newblock {\em European Journal of Operational Research}, 234(3):597--609,
  2014.

\bibitem{lopez2016irace122}
Manuel L{\'o}pez-Ib{\'a}{\~n}ez, J{\'e}r{\'e}mie Dubois-Lacoste,
  Leslie~P{\'e}rez C{\'a}ceres, Mauro Birattari, and Thomas St{\"u}tzle.
\newblock The irace package: Iterated racing for automatic algorithm
  configuration.
\newblock {\em Operations Research Perspectives}, 3:43--58, 2016.

\bibitem{yang2017adaptive187}
Qiang Yang, Wei-Neng Chen, Zhengtao Yu, Tianlong Gu, Yun Li, Huaxiang Zhang,
  and Jun Zhang.
\newblock Adaptive multimodal continuous ant colony optimization.
\newblock {\em IEEE Transactions on Evolutionary Computation}, 21(2):191--205,
  2017.

\bibitem{oliveira2017analysis176}
Sabrina Oliveira, Mohamed~Saifullah Hussin, Andrea Roli, Marco Dorigo, and
  Thomas St{\"u}tzle.
\newblock Analysis of the population-based ant colony optimization algorithm
  for the tsp and the qap.
\newblock In {\em Evolutionary Computation (CEC), 2017 IEEE Congress on}, pages
  1734--1741. IEEE, 2017.

\bibitem{stutzle2018iterated185}
Thomas St{\"u}tzle and Rub{\'e}n Ruiz.
\newblock Iterated greedy.
\newblock {\em Handbook of Heuristics}, pages 547--577, 2018.

\bibitem{hara2010ant182}
Akira Hara, Syuhei Matsushima, Takumi Ichimura, and Tetsuyuki Takahama.
\newblock Ant colony optimization using exploratory ants for constructing
  partial solutions.
\newblock In {\em Evolutionary Computation (CEC), 2010 IEEE Congress on}, pages
  1--7. IEEE, 2010.

\bibitem{blum2011hybrid139}
Christian Blum, Jakob Puchinger, G{\"u}nther~R Raidl, and Andrea Roli.
\newblock Hybrid metaheuristics in combinatorial optimization: A survey.
\newblock {\em Applied Soft Computing}, 11(6):4135--4151, 2011.

\bibitem{solnon2013ant105}
Christine Solnon.
\newblock {\em Ant colony optimization and constraint programming}.
\newblock John Wiley \& Sons, 2013.

\bibitem{massen2012pheromone128}
Florence Massen, Yves Deville, and Pascal Van~Hentenryck.
\newblock Pheromone-based heuristic column generation for vehicle routing
  problems with black box feasibility.
\newblock In {\em International Conference on Integration of Artificial
  Intelligence (AI) and Operations Research (OR) Techniques in Constraint
  Programming}, pages 260--274. Springer, 2012.

\bibitem{massen2013experimental129}
Florence Massen, Manuel L{\'o}pez-Ib{\'a}nez, Thomas St{\"u}tzle, and Yves
  Deville.
\newblock Experimental analysis of pheromone-based heuristic column generation
  using irace.
\newblock In {\em International Workshop on Hybrid Metaheuristics}, pages
  92--106. Springer, 2013.

\bibitem{merkle2002modeling134}
Daniel Merkle and Martin Middendorf.
\newblock Modeling the dynamics of ant colony optimization.
\newblock {\em Evolutionary Computation}, 10(3):235--262, 2002.

\bibitem{uchida2012efficient123}
Akihiro Uchida, Yasuaki Ito, and Koji Nakano.
\newblock An efficient gpu implementation of ant colony optimization for the
  traveling salesman problem.
\newblock In {\em Networking and computing (icnc), 2012 third international
  conference on}, pages 94--102. IEEE, 2012.

\bibitem{twomey2010analysis184}
Colin Twomey, Thomas St{\"u}tzle, Marco Dorigo, Max Manfrin, and Mauro
  Birattari.
\newblock An analysis of communication policies for homogeneous multi-colony
  aco algorithms.
\newblock {\em Information Sciences}, 180(12):2390--2404, 2010.

\bibitem{guerrero2014comparative90}
Gin{\'e}s~D Guerrero, Jos{\'e}~M Cecilia, Antonio Llanes, Jos{\'e}~M
  Garc{\'\i}a, Martyn Amos, and Manuel Ujald{\'o}n.
\newblock Comparative evaluation of platforms for parallel ant colony
  optimization.
\newblock {\em The Journal of Supercomputing}, 69(1):318--329, 2014.

\bibitem{cecilia2013enhancing35}
Jos{\'e}~M Cecilia, Jos{\'e}~M Garc{\'\i}a, Andy Nisbet, Martyn Amos, and
  Manuel Ujald{\'o}n.
\newblock Enhancing data parallelism for ant colony optimization on gpus.
\newblock {\em Journal of Parallel and Distributed Computing}, 73(1):42--51,
  2013.

\bibitem{dawson2013improving43}
Laurence Dawson and Iain Stewart.
\newblock Improving ant colony optimization performance on the gpu using cuda.
\newblock In {\em Evolutionary Computation (CEC), 2013 IEEE Congress on}, pages
  1901--1908. IEEE, 2013.

\bibitem{delevacq2013parallel46}
Audrey Del{\'e}Vacq, Pierre Delisle, Marc Gravel, and Micha{\"e}l Krajecki.
\newblock Parallel ant colony optimization on graphics processing units.
\newblock {\em Journal of Parallel and Distributed Computing}, 73(1):52--61,
  2013.

\end{thebibliography}

\documentclass[runningheads]{llncs}
\usepackage{graphicx}
\usepackage{listings}
\usepackage{algorithm2e}
\begin{document}
\title{Evolution of Ant Colony Optimization Algorithm --- A Brief Literature Review}
%
%
\author{Aleem Akhtar}

	\institute{SEECS-NUST, ISLAMABAD \\
		\email{aleem.akhtar@seecs.edu.pk} \\
		 }
%
%
%
\maketitle              
\begin{abstract}
Ant Colony Optimzation (ACO) is a metaheuristic proposed by Marco Dorigo in 1991 based on behavior of biological ants. Pheromone laying and selection of shortest route with the help of pheromone inspired development of first ACO algorithm. Since, presentation of first such algorithm, many researchers have worked and published their research in this field. Though initial results were not so promising but recent developments have made this metaheuristic a significant algorithm in Swarm Intelligence. This research presents a brief overview of recent developments carried out in ACO algorithms in terms of both applications and algorithmic developments. For application developments, multi-objective optimization, continuous optimization and time-varying NP-hard problems have been presented. While to review articles based on algorithmic development, hybridization and parallel architectures have been investigated.

\keywords{ACO Algorithm \and ACO  Metaheuristic \and Evolution \and Literature Review.}
\end{abstract}
\section{Introduction}
Ant Colony Optimization or simply ACO is a meateuristic that is used to solve optimization problems that are complex combinatorial ~\cite{dorigo1999ant,Dorigo1999}. Basic algorithm of ACO is based on laying of pheromone trail inspired by biological ants’ behavior, which use pheromones as medium of communication. In a similar way to the biological example, simple agents – artificial ants – communicating with each other indirectly using artificial pheromones within a colony is base of ACO. ACO is based on indirect communication within a colony of simple agents, called ants (artificial), induced by (artificial) pheromone bands. Traces of pheromones work as distributed and digital information in ACO which ants use to build solutions to the problem to be solved and to adapt to it during the execution of the algorithm as a reflection of their search experience which is a major difference between ACO algorithm and other construction heuristics.
Ant System (AS) ~\cite{dorigo1992optimization,dorigo1991ants,dorigo1996ant69} was one of the starting example of such algorithm that was put forward using an eminent application example – Travelling Salesman Problem (TSP) ~\cite{applegate2006traveling6,reinelt1994traveling155}. Though Ant System presented promising results but it was still not enough to compete with the most advanced TSP algorithms. Still, it played an important role to encourage further exploration of the two algorithmic variants that provided significantly better computing power and applications. In fact, quite astonishing results are achieved in significantly large number of applications using this algorithm. Problems like DNA sequencing ~\cite{blum2008ant27}, scheduling ~\cite{blum2005beam20}, protein-ligand docking ~\cite{korb2007ant107}, assembly line balancing ~\cite{blum2008beam21}, sequential ordering ~\cite{gambardella2000ant84}, packet-switched routing ~\cite{di1998antnet52} and 2D-HP protein folding ~\cite{shmygelska2005ant160} are some of example applications of ACO.  A common framework is provided by ACO metaheuristic for algorithmic variations and prevailing applications ~\cite{dorigo1999ant}. ACO metaheuristic following algorithms are called ACO algorithms. 
Hundreds of researchers have been attracted towards with ACO field since early version of ACO algorithm was proposed in 1991. ACO is now considered as a heavily developed metaheuristic and there exists a great number of both theoretical and experimental research results by now. ACO importance can be demonstrated by ~\cite{doerner2009special57}:
\begin{enumerate}
\item	IEEE Swarm Intelligence Symposium series
\item	Biannual International Conference on Ant Colony Optimization and Swarm Intelligence (ANTS)
\item	Several conferences on evolutionary algorithms and metaheuristics
\item	Special journal issues
\item	Dedicated website and mailing list for ACO (www.aco-metaheuristic.org)
\end{enumerate}

The rest of this article is divided into four sections. The first section, presents a brief background of ACO algorithm followed by initial developments performed in ACO algorithm in the second section. The third section, gives a detailed overview of recent developments (after 2010) in terms of application development, algorithmic development and parallel implementation. The fourth and last section concludes this article. 

\section{Historical Background}
Ant System (AS) was preliminary proposed ACO algorithm that was applied to relatively lesser TSP instances not more than 80 cities. Performance matching some other common heuristics like evolutionary computation ~\cite{dorigo1992optimization,dorigo1996ant69} was achieved by initial AS yet it was not enough to match or compete performance given by other advanced algorithms designed for TSP. As a result, focus of research within the ACO has been shifted towards ACO algorithms that work better than AS, for example when applied to the TSP. The rest of this section, provides introduction of biological metaphor that was inspiration for both AS and ACO. 
\subsection{Biological Analogy}
In most ant species, a chemical known as pheromone is deposited on ground by ants while walking which is smelled by other ants ~\cite{deneubourg1990self48}. From nest to food source, a path is marked through a trail created due to pheromones deposited by ants. Other ants in search of food, use the pheromone trail (path) to locate the food. Many ant species are proficient in determining shortest trail from nest to food location by exploiting pheromone trails. Deneubourg et al ~\cite{deneubourg1990self48}, created a controlled experimental environment to study and research ants behavior in pheromone laying and determination of shortest path to food. For this purpose, they used a variable length double bridge to connect food source and nest of ants. They ran multiple experiments with one longer bridge and other shorter bridge. Results showed that, in the beginning of each experiment, ants started a random motion towards the food in the arena but shorter path was used by ants in the end. Results presented by Deneubourg et al ~\cite{deneubourg1990self48}, can be explained as follows. In the start, no ant travelled the environment therefore there was pheromone present on the ground which gives same probability of any route selection without any preference. It can be said that, on average nearly half number of ants chose each path in the beginning. However, those ants which selected shorter path reached the food source and travelled back to nest earlier than other set of ants which selected longer path. Now, to again travel back towards food source, ants decide to choose one path and due to presence of higher pheromone level on shorter path prejudice their decision. In the end, pheromone level on shorter path start to add up faster causing nearly entire colony of ants to use that path. 
In Ant Systems and different variants of ACO algorithms, same phenomenon is applied. Ants are replaced by artificial ants (agents). Artificial pheromone trails are used to represent pheromone trails and graph is used to represent double bridge. Extra capabilities to implement restrictions and retracing of solution without errors are given to artificial ants to solve real life problems. Quantity of pheromone deposited by artificial ants is set at proportional to quality of produced solution – a behavior similar to biological ants ~\cite{beckers1993modulation10}.

\section{Initial Developments}
Marco Dorigo was first person to propose and publish first ACO algorithm – Ant System (AS) – in a set of three algorithms namely ant-quantity, ant-density, and ant-cycle as part of his doctoral thesis ~\cite{dorigo1992optimization}. Few years later, these three algorithm appeared first as technical report ~\cite{dorigo1991ants} in the IEE Transactions on Man, Cybernetics and Systems ~\cite{dorigo1996ant69}. Difference between three algorithms was that in ant-quantity and ant-density, pheromone was updated right after ants move from one city to another, while in ant-cycle, pheromone deposition was updated once all ants had built the path and tour quality was used as a function for level of pheromone update. Better performance presented by ant-cycle caused research in other two algorithms to stop and ant-cycle was used to present Ant System. Initial version of AS, presented encouraging results but were not enough to compete with other well-established algorithms. However, these results were encouraging enough to stimulate exploration of research in this field.
\subsection{ACO Algorithm}
Basic ACO algorithm is given in Figure \ref{ACOAlgo} and consists of four main stages. Each stage is explained as follow:
\begin {itemize}
\item \textbf{Initialization}: This is first step, in which all parameters and pheromones variables are set
\item \textbf{Construct Ant Solutions}: After initialization, a set of ants build solution to the problem being solved using pheromones values and other information
\item \textbf{Local Search}: This stage is optional and it includes improvement of constructed solution by ants
\item \textbf{Global Pheromone Update}: This is last stage, and it includes update in pheromone variables based on search experience reflected by ants.
\end {itemize}

\begin{algorithm}[H]
	\textbf{procedure} \textit{Start of ACO Algorithm}\\
	initialization\;
	\While{(Iterate until end criteria is reached)}{
		BuildAntSolutions\;
		LocalSearchMethod  \% optional \;
		PheromonesGlobalUpdate\;
	}
\textbf{end} \textit{End of ACO algorithm}
	\caption{ACO Algorithm}
	\label{ACOAlgo}
\end{algorithm}

\subsection{Earlier Variants of ACO}
There have been numerous updates and improvements made in ACO algorithms which are presented and published by many researchers in the literature. A brief overview on each variant can take a lot time to report therefore this article only focuses on details of development done in ACO algorithm after year 2010. Earlier variants of ACO algorithms with proposal year and main references are presented in Table \ref{ealierDev}.

\begin{table}
	\centering
	\caption{Earlier Developments in ACO Algorithm}
	\label{ealierDev}
\begin{tabular}{|c|c|c|}
	\hline 
\textbf{	ACO Algorithm} & \textbf{Year} & \textbf{Main Reference} \\ 
	\hline 
	Ant System & 1991 & ~\cite{dorigo1992optimization,dorigo1991ants,dorigo1996ant69} \\ 
	\hline 
	Elitist AS & 1992 & ~\cite{dorigo1992optimization,dorigo1991ants,dorigo1996ant69} \\ 
	\hline 
	Ant-Q & 1995 &  ~\cite{gambardella1995ant82} \\ 
	\hline 
	Ant Colony System & 1996 &  ~\cite{dorrigo1997ant64} \\ 
	\hline 
	Max-Min Ant System & 1996 & ~\cite{stutzle1996improving174,stutzle1997max175} \\ 
	\hline 
	Rank-based AS & 1997 & ~\cite{bullnheimer1997new31} \\ 
	\hline 
	ANTS & 1998 & ~\cite{maniezzo1999exact124} \\ 
	\hline 
	Best-worst AS & 2000 & ~\cite{cordon2002analysis38,cordon2000new39} \\ 
	\hline 
	Population based ACO & 2002 & ~\cite{guntsch2002population92} \\ 
	\hline 
	Beam-ACO & 2004 & ~\cite{blum2004theoretical19,blum2005beam20} \\ 
	\hline 
	Hyper Cube – ACO & 2004 & ~\cite{blum2001hcHCACO} \\ 
	\hline 
\end{tabular}
\end{table} 

\section{Developments}
As mentioned earlier, covering updates and improvements in ACO algorithm from initial proposed system will take a lot of time, therefore, this review article covers recent evolutions in ACO algorithms after year 2010. In this section, latest research updates in ACO are presented. This section is divided into three sub sections. The first subsection, outlines application of Ant Colony Optimization algorithms to non-standard problems followed by ACO developments from algorithmic point of view in the second subsection. The third and the final subsection presents improvements in ACO algorithm with parallel implementations.
\subsection{ACO Non-standard Applications}
This subsection presents a review of ACO application to problems involving factors such as time-varying data, multi-objective function and continuous optimization.
\subsubsection{NP-hard Problems}
Dynamic problems where characteristics change during runtime such as network routing are main applications in which ACO algorithm have been significantly successful. Classical NP-hard problems with dynamic variants is one of main area where ACO algorithms have also been applied. One such example is dynamic version of Travelling Salesman Problem (TSP) where cities may disappear or appear or distance between cities may change. Mavrovouniotis et al ~\cite{mavrovouniotis2017ant133} presented a more recent work in this field that targeted towards improvement in performance of ACO on dynamic problems by explicitly using local search algorithms. Mavrovouniotis and Yang in 2015 ~\cite{mavrovouniotis2015ant131} reported ACO algorithm application to dynamic vehicle routing problems with better results on both real-world instances and academic instances. Mavrovouniotis et al ~\cite{mavrovouniotis2017survey132} published another review article on dynamic optimization problems with swarm intelligence algorithms such as ACO. 
\subsubsection{Multi-Objective Optimization}
In real-world applications, multiple solutions are assessed as a function of several, mostly differing objectives. Different approaches are used to handle these problems. One approach is to order objectives according to their importance as reported by Veen et al ~\cite{van2013ant85} as part of their research. They presented a two-colony Ant Colony System (ACS) algorithm – MACS-VRPTW – to solve routing problem of vehicles with time windows. López-Ibáñez and Stützle proposed a different approach to develop multi-objective ACO algorithms ~\cite{lopez2012experimental119}. Different existing methods to manage multi-objective problems were analyzed by them and a generic multi-objective ACO version (MO-ACO) is proposed. This version can be used to instantiate other approaches as well as new variations can also be developed. By exploring the design space created by MOACO algorithms with the new methodology of creating multi-objective optimizers, new MOACO algorithms were automatically generated that outperformed all the earlier proposed ACO algorithms in this field ~\cite{lopez2012automatic118}. In more recent research, Falcón-Cardona and Coello Coello further extended the framework presented by Stützle et al to present a novel approach for multi-objective problems with ACO algorithm variant called iMOACO-RR ~\cite{falcon2017new77}.
\subsubsection{Continuous Optimization}
Initially ACO was applied to combinatorial problems but later on it was adapted for continuous optimization problems. Variables’ discretization of real-valued domain is simplest methodology to apply ACO to continuous problems. Implementation of ACO using this approach has been applied to Protein-ligand docking problem. Socha and Dorigo first presented an extended version of ACO named as ACOR algorithm where Gaussian kernel functions are explicitly used in place of probability density functions. Liao et al presented a refined form of ACOR algorithm using growing population size and incorporating dominant local search algorithms ~\cite{liao2011incremental113}. Kumar et al ~\cite{kumar2015enhancing109} and Guo et al ~\cite{guo2012antNew}, later on presented refined forms of same algorithm. Liao et al ~\cite{liao2014unified114} presented an integrated structure for ACO applications to continuous optimization. This version of ACO allows selection of particular algorithm components to instantiate different variants of ACOR. Unified structure was able to generate ACO algorithms using Irace – an automatic algorithm setting tool, that outperformed all the earlier versions presented in the literature ~\cite{lopez2016irace122}. Yang et al ~\cite{yang2017adaptive187}, also presented an extended version of ACOR for multi-modal optimization.
\subsection{Developments in Algorithms}
This sub section presents review of research articles that mainly focused on development of ACO variants to improve performance. This includes ACO hybridization with other algorithms and inter-programming techniques. 
\subsubsection{ACO Combination with Other Metaheuristics}
To fine-tune solutions built by ants, local improvement heuristics are combined with ACO as most common technique. For this purpose, repetitive improvement algorithms are mainly used but many other articles have reported use of other metaheuristic algorithms for improvement. Oliveira et al ~\cite{oliveira2017analysis176} presented use of tabu search for quadratic assignment problem in order to improve solution constructed by ants. Many other hybridization techniques have been proposed such as letting ants to build a solution from partially present solution as reported by Stutzle et al ~\cite{stutzle2018iterated185} in their article in the form of iterated ants. Another approach presented by Hara et al ~\cite{hara2010ant182} is to use other complete solutions to take out partial solutions. Using former approach, fasten the process of solution construction as well allows exploitation of good parts of solution directly. Results presented by researchers using this approach showed that this approach is quite effective in the absence of other useful local search methods. 
\subsubsection{ACO Combination with Inter-Programming Techniques}
Integration of ACO with programming techniques is very useful for highly constrained problems for which finding a feasible solution is not so easy. Blum et al ~\cite{blum2011hybrid139} presented an approach that require integration of ants’ solutions construction with constraint propagation mechanism to determine if certain solution will be feasible or not. This approach have shown good results for scheduling problems that are highly constrained.Solnon et al ~\cite{solnon2013ant105}, recently published his research that focuses on integration of constraint solver into ACO algorithm. Column generation technique is integrated with ACO to solve vehicle routing problems that include black-box achievability constraints is presented by Massen et al ~\cite{massen2012pheromone128}. This research was further explored and tested on various benchmarks with improved results in extended version of research article by Massen et al ~\cite{massen2013experimental129}.
\subsection{Parallel Implementations}
The nature of the ACO algorithms allows parallelization to be applied to both population and data domain. Other algorithms based on population using parallelization schemes can be upgraded to ACO. Course grained and fine grained techniques are used to as classification of older parallelization strategies. Fine parallelization is characterized by the fact that very few people are assigned to a single processor and that information is exchanged frequently between processors. On the other hand, coarse-grained approaches use bigger population sometime entire population to be attributed to individual processor and the exchange of information is rather rare. Merkle et al ~\cite{merkle2002modeling134} provides a brief overview of both approaches.
Fine-grained parallelism schemes were explored very early, when shared memory models and multicore CPUs were not common or unavailable. Initial fine-grained strategy were investigated with parallel AS versions for the TSP on the CM-2 Connection Machine by assigning each Ant to a single processing unit. However, results presented by experiments using this approach reported negative values due to communication overhead between ants for modification of pheromone trails. On other hand, research have showed that coarse-grained strategies are much more effective for ACO. For implementing course-grained parallel technique to ACO, p sub-colonies are executed in parallel depending on p available processors. Uchida et al ~\cite{uchida2012efficient123} used this technique to present results that proved running p independent sub-colonies an effective technique with further improvements achieved by improving exchange of information among sub-colonies. These improvements were studied and presented in detail by Twomey et al ~\cite{twomey2010analysis184} in their research. Presence of shared memory models and multicore CPU architecture allows introduction of thread level parallelism into ACO algorithm that can further improve speed of ACO algorithm. Guerrero et al ~\cite{guerrero2014comparative90} evaluated parallel versions of ACO algorithms on various platforms while most recent trends have seen implementation of ACO algorithms on GPUs to speed-up, as reported by ~\cite{cecilia2013enhancing35,dawson2013improving43,delevacq2013parallel46}.  
\section{Conclusion}
Since its introduction in 1991 by Dorigo, ACO algorithm has become one of most widely used metaheuristic for solving combinatorial problems. Started as Ant System (AS), initial versions of ACO were not promising enough to compete with other established algorithms but results were encouraging enough to open up the gates of research in this field. Since then, many researchers have exploited basic version of ACO to work on bringing updates in it for achieving better results. This research focused on outlining more recent developments carried out in ACO algorithms in terms of both applications and algorithmic developments. For non-standard applications of ACO algorithms, multi-objective optimization, continuous optimization and time-varying NP-hard problems were main focus of recent developments. While to improve, performance of ACO algorithm, it has been integrated with other metaheuristics and inter-programming techniques. Hybridization of ACO algorithms have clearly showed a lot of improvement in results for various problems. Recent trends have also seen implementation of ACO algorithms with parallel versions. Availability of multicore CPU architectures and GPUs have made it possible to develop improved parallel versions of ACO algorithms.  

\bibliographystyle{unsrt}


\bibliography{acobib}

\end{document}